The y in the the price level. percentage change in the price level as measared by that price index Contractions are periods when the real GDP grows slower than the tread occurring. worker health may be advensely affected worker quality may be adversely affectod unemployment may increase.
Answer #1

17) Quantity theory is given by P x Y = M x V where Y is the real GDP or real Output. (total output in the economy, first option)

18) Inflation rate = (P1/P0)^(1/n) - 1 = (140/120)^(1/1) - 1 = 16.67%. Select 17% (second option)

19) Contraction and expansions are phases of business cycles. Contractions occur when economic growth is low and this is followed by trough (third option)

20) Unemployment rises because fewer workers will be hired